OpenStudy (lgbasallote):

A vector A of magnitude 6 and another vector B have a resultant magnitude 12. The vector B: a. must have a magnitude of at least 6 but no more than 18 b. may have a magnitude of 20 c. cannot have a magnitude greater than 12 d. must be perpendicular to A the formula for resultant vector is \[R = \sqrt{A^2 + B^2}\]isn't it?

OpenStudy (anonymous):

that only works if the triangle of vector A, vector B, & vector R form a Right Triangle

OpenStudy (anonymous):

think about using the rules for making triangles to assist you in this problem

OpenStudy (lgbasallote):

don't all vectors form right triangles?

OpenStudy (anonymous):

no, i will draw one that doesn't

OpenStudy (anonymous):



OpenStudy (lgbasallote):

i guess that's a point...so what's the general idea then?

OpenStudy (anonymous):

if the formula you have written words than your answer is d

OpenStudy (anonymous):

The sum of two sides of a triangle cannot be less than the remaining side

OpenStudy (lgbasallote):

i meant with vectors

OpenStudy (ash2326):

if both vectors are in the same direction then the second could be of magnitude 6 so 6+6=12 if in a perpendicular direction then \[\sqrt {6^2+x^2}=12\] or \[x^2=108\] \[x=10.39\] @lgbasallote do you get the idea?

OpenStudy (lgbasallote):

that's the idea there too?

OpenStudy (lgbasallote):

what's that 6 + 6 = 12 for?

OpenStudy (ash2326):

if they are in the same direction

OpenStudy (lgbasallote):

you lost me there..i do know that the third side could be 10.38...

OpenStudy (lgbasallote):

anyway...why is B perpendicular to A again @juantweaver ?

OpenStudy (ash2326):

@lgbasallote sorry I made a mistake, let me see it again

ganeshie8 (ganeshie8):